# ReviewEval: An Evaluation Framework for AI-Generated Reviews

### v2

@@ Abstract
The escalating volume of academic research, coupled with a shortage of qualified reviewers, necessitates innovative approaches to peer review. While large language model (LLMs) offer potential for automating this process, their current limitations include superficial critiques, hallucinations, and a lack of actionable insights. This research addresses these challenges by introducing a comprehensive evaluation framework for AI-generated reviews, that measures alignment with human evaluations, verifies factual accuracy, assesses analytical depth, and identifies actionable insights. We also propose a novel alignment mechanism that tailors LLM-generated reviews to the unique evaluation priorities of individual conferences and journals. To enhance the quality of these reviews, we introduce a self-refinement loop that iteratively optimizes the LLM’s review prompts. Our framework establishes standardized metrics for evaluating AI-based review systems, thereby bolstering the reliability of AI-generated reviews in academic research.
@@
Chhavi Kirtani*1, Madhav Krishan Garg*1, Tejash Prasad*1, Tanmay Singhal*1, Murari Mandal2, Dhruv Kumar1, 3, 1IIIT Delhi, 2KIIT Bhubaneswar, 3BITS Pilani, Correspondence: dhruv.kumar@pilani.bits-pilani.ac.in *Equal Contributions
@@ 1 Introduction
The rapid growth of academic research, coupled with a shortage of qualified reviewers, has created an urgent need for scalable and high-quality peer review processes Petrescu and Krishen (2022); Schulz et al. (2022); Checco et al. (2021). Traditional peer review methods are under mounting pressure from the exponentially growing number of submissions, particularly in fields like artificial intelligence, machine learning, and computer vision. This has led to a growing interest in leveraging large language models (LLMs) to automate and enhance various aspects of the peer review process Robertson (2023); Liu and Shah (2023).
LLMs have shown remarkable potential in automating various natural language processing tasks, such as summarization, translation, and question-answering. However, their effectiveness in serving as reliable and consistent paper reviewers remains a significant challenge. The academic community is already experimenting with AI-assisted reviews, as evidenced by reports that 15.8% of reviews for ICLR 2024 were generated with AI assistance Latona et al. (2024). While this demonstrates the growing adoption of LLMs in peer review, concerns have been raised regarding their impact on the reliability and fairness of the review process. Specifically, papers reviewed by AI have been perceived to gain an unfair advantage, leading to questions about the integrity of such evaluations. Consequently, research into robust automated review generation systems is crucial, necessitating rigorous evaluation of AI generated reviews to address key challenges. Zhou et al. (2024) provide a comprehensive analysis of the use of commercial models, such as GPT-3.5 and GPT-4 Achiam et al. (2023), as research paper reviewers. Their findings highlight key limitations, including the potential for mistakes due to either model hallucinations or an incomplete understanding of the material, as well as the inability to provide critical feedback comparable to human reviewers. Based on our preliminary experiments of using GPT-4 to generate reviews for research papers, we identified additional limitations in AI-generated reviews, including a lack of actionable insights and limited analytical depth, often characterized by generic and vague feedback. These shortcomings stem from the inherent tendency of LLMs to generate superficial reviews.
@@
Existing research on evaluation metrics for AI-generated research paper reviews remains limited. For instance, D’Arcy et al. (2024) proposed an automated metric to evaluate approximate matches between AI-generated and human-written review comments using GPT-4 Achiam et al. (2023). Although their method iteratively employs GPT-4 to extract approximate matches and mitigate inconsistencies, its complete reliance on GPT-4 renders the evaluation process a black box, thereby limiting transparency and raising concerns about reliability. Similarly, Zhou et al. (2024) investigated the aspect coverage and similarity between AI and human reviews through a blend of automatic metrics and manual analysis. Their work leveraged the ASAP dataset Yuan et al. (2022), which categorizes review sentences into predefined aspects—including summary, motivation, originality, soundness, substance, replicability, meaningful comparison, and clarity, to align AI and human reviews. However, beyond this AI-human comparison, their approach overlooks other critical dimensions where AI reviews may underperform, as highlighted in Figure 1.
@@
Based on our analysis of the limitations in current AI-generated reviews and the gaps in existing evaluation metrics, we propose a comprehensive evaluation framework designed to assess the quality of AI-generated research paper reviews. Our framework targets four key dimensions (see Figure 1): ❶ Comparison with Human Reviews: Evaluates topic coverage and semantic similarity to measure the alignment between AI-generated and human-written feedback. ❷ Factual Accuracy: Detects factual errors, including misinterpretations, incorrect claims, and hallucinated information. ❸ Analytical Depth: Assesses whether the AI’s critique transcends generic commentary to offer in-depth, meaningful engagement with the research. ❹ Actionable Insights: Measures the ability of the AI to provide specific, constructive suggestions for improving the paper.
@@
Recognizing that major conferences and journals have distinct reviewing priorities, a one-size-fits-all approach to AI-driven reviews is insufficient. Recent studies Bauchner and Rivara (2024); Biswas (2024) underscore the growing importance of aligning reviews with conference-specific evaluation criteria—especially as many venues now require adherence to detailed reporting guidelines. To address this, we introduce a conference-specific AI reviewer that dynamically adapts its review strategy to meet the unique criteria of each target venue.
Furthermore, inspired by the self-refinement approach of Madaan et al. (2023), our system incorporates a supervisor model that iteratively critiques and refines the instructional prompts guiding the review process. This self-refinement loop promotes deeper analytical assessments and mitigates the tendency of AI-generated reviews to offer only superficial feedback. In summary, the paper makes the following contributions:
  1. 1.
    Develop a comprehensive evaluation framework for LLM-based reviewing systems across five dimensions: (i) alignment with human reviews, (ii) factual accuracy, (iii) analytical depth, (iv) actionable insights, and (v) adherence to Reviewer guidelines.
  2. 2.
    Create an LLM-based reviewer that dynamically aligns with the evaluation criteria of specific conferences/journals and continuously improves its reviewing strategy through an iterative refinement loop.
@@ 2 Related Work
AI-based scientific discovery. Early efforts on automating scientific discovery include expert systems from 1970s such as DENDRAL Buchanan and Feigenbaum (1981) and the automated mathematician Lenat (1977), which focused on constrained problem spaces like organic chemistry and theorem proving. More recently, LLMs and machine learning techniques have been used to extend automated research beyond structured domains. Notable contributions include AutoML that optimize hyperparameters and architectures Hutter et al. (2019); He et al. (2021) and AI-driven discovery in materials science and synthetic biology Merchant et al. (2023); Hayes et al. (2024). However, these methods remain largely dependent on human-defined search spaces and predefined evaluation metrics, limiting their potential for open-ended discovery. Recent works Lu et al. (2024) aim to automate the entire research cycle, encompassing ideation, experimentation, manuscript generation, and peer review, thus pushing the boundaries of AI-driven scientific discovery.
Figure 1: We show some examples of the challenges and limitations of AI generated research paper reviews
@@
Lu et al. (2024) employ LLMs to autonomously conduct the research pipeline, including peer review. It follows a structured three-stage review process: paper understanding, criterion-based evaluation (aligned with NeurIPS and ICLR guidelines), and final synthesis: assigning scores to key aspects like novelty, clarity, and significance. Evaluation shows its reviews closely match human meta-reviews. MARG D’Arcy et al. (2024) introduces a multi-agent framework where worker agents review sections, expert agents assess specific aspects, and a leader agent synthesizes feedback. Using BERTScore Zhang* et al. (2020) and GPT-4-based evaluation, MARG-S improves feedback quality, reducing generic comments and increasing helpful feedback per paper. These studies highlight the AI’s potential to enhance peer review through structured automation and multi-agent collaboration.
Evaluation framework for AI-based peer-review. There has been limited research on developing evaluation frameworks for evaluating the quality of LLM generated paper reviews. Zhou et al. (2024) evaluated GPT models for research paper reviewing across 3 tasks: aspect score prediction, review generation, and review-revision MCQ answering. The study developed evaluation framework comprising of aspect coverage (originality, soundness, substance, replicability, etc.), ROUGE (lexical overlap), BERTScore (semantic similarity), and BLANC (informativeness), alongside manual analysis. Results showed LLMs overemphasized positive feedback, lacked critical depth, and neglected substance and clarity, despite high lexical similarity to human reviews. D’Arcy et al. (2024) introduced an automated evaluation framework for AI-generated reviews, quantifying similarity to human reviews via recall, precision, and Jaccard index. Recall measures the fraction of real-reviewer comments with at least one AI match, precision quantifies AI comments aligned with human reviews, and Jaccard index evaluates the intersection-over-union of aligned comments.
@@
Table 1: Existing evaluation metrics focus on AI and human review similarity but overlook other key metrics. Our framework addresses these gaps with a more interpretable metrics.
@@ 3 ReviewEval
We describe the development of our evaluation framework and the LLM-based research paper reviewer. Each review is evaluated on several key parameters to assess the overall quality of the generated feedback. To ensure consistency and reliability, all evaluations for a given metric were performed by LLMs of the same specification and version, thereby maintaining inter-rater reliability and ensuring robust, unbiased comparisons. Our experiments utilized 16 papers and their corresponding expert reviews (scraped from OpenReview.net), alongside AI-generated reviews. The overview of the proposed framework is presented in Figure 2.
@@ 3.1 Comparison with Expert Reviews
We compare the reviews generated by the LLM based reviewer with expert reviews from OpenReview.net. Our primary goal is to gauge how well the AI system replicate or complement expert-level critique. The evaluation is conducted along the following dimensions:
Semantic similarity. To assess the alignment between AI-generated and expert reviews, we embed each review $R$ into a vector space using the OpenAI embedding model. The semantic similarity between an AI-generated review $R_{\text{AI}}$ and an expert review $R_{\text{Expt}}$ is measured using cosine similarity:
@@
Figure 2: Schematic overview of the ReviewEval framework. Given a paper and the associated conference or journal guidelines, the system generates AI-based reviews and evaluates them along multiple dimensions.
Topic modeling. We evaluate topic coverage to determine how comprehensively AI-generated reviews address the breadth of topics present in expert reviews. Our approach comprises three steps: ❶ Topic extraction: Each review $R$ (either AI-generated or expert) is decomposed into a set of topics: $T_{R}=\{t_{1},t_{2},\dots,t_{n}\},$ where each topic $t_{i}$ is represented by a sentence that captures its core content and context. ❷ Topic similarity: Let $T_{\text{AI}}=\{t_{1},t_{2},\dots,t_{m}\}$ and $T_{\text{Expt}}=\{t^{\prime}_{1},t^{\prime}_{2},\dots,t^{\prime}_{n}\}$ denote the topics extracted from the AI and expert reviews, respectively. We define a topic similarity function $\text{TS}(t_{i},t^{\prime}_{j})$ that an LLM assigns on a discrete scale:
$\begin{split}\text{TS}(t_{i},t^{\prime}_{j})=3\cdot\mathbb{I}\{t_{i}\sim_{\text{strong}}t^{\prime}_{j}\}\\ +2\cdot\mathbb{I}\{t_{i}\sim_{\text{moderate}}t^{\prime}_{j}\}\\ +1\cdot\mathbb{I}\{t_{i}\sim_{\text{weak}}t^{\prime}_{j}\},\end{split}$ (2)
where $\mathbb{I}$ is the indicator function, $t_{i}\sim_{\text{strong}}t^{\prime}_{j}$, $t_{i}\sim_{\text{moderate}}t^{\prime}_{j}$ , $t_{i}\sim_{\text{weak}}t^{\prime}_{j}$ denote substantial, moderate, and minimal overlap in concepts, respectively. All the conditions are mutually exclusive. We set a similarity threshold $\tau=2$ so that topics with $\text{TS}(t_{i},t^{\prime}_{j})\geq\tau$ are considered aligned. ❸ Coverage ratio: For each AI-generated review, we construct a topic similarity matrix $S$ where each element $S[i,j]=\text{TS}(t_{i},t^{\prime}_{j})$ represents the similarity between topic $t_{i}$ from $T_{\text{AI}}$ and topic $t^{\prime}_{j}$ from $T_{\text{Expt}}$. The topic coverage ratio is defined as:
@@
3.2 Evaluating Factual Correctness of Reviews
To address the hallucinations and factual inaccuracies of LLM generated reviews, we propose an automated pipeline that validates the factual correctness of LLM-generated reviews by simulating the conference rebuttal process. Our pipeline emulates the traditional rebuttal workflow, where authors clarify or counter reviewer claims using evidence from their work. By automating both the question generation and rebuttal phases, our system produces a robust factual correctness evaluation. The pipeline consists of the following steps:
Step 1: Transforming reviews into structured questions. Each LLM-generated review $R$ is transformed into a structured question $Q$ that encapsulates the central claim or critique. For example, consider the following review from the PeerRead dataset Kang et al. (2018) for the paper "Augmenting Negative Representations for Continual Self-Supervised Learning" Cha et al. :
Review ($R$):Augmentation represents a crucial area of exploration in self-supervised learning. Given that the authors classify their method as a form of augmentation, it becomes essential to engage in comparisons and discussions with existing augmentation methods.
This review is converted into the corresponding question:
Generated question ($Q$):Has the paper engaged in comparisons and discussions with existing augmentation methods, given that the authors classify their method as a form of augmentation?
Step 2: Decomposing questions into sub-questions. The question $Q$ is then decomposed into a set of sub-questions $\{q_{1},q_{2},\ldots,q_{n}\}$ using a dedicated query decomposition engine. This decomposition enables a fine-grained analysis by isolating distinct components of the original question.
Step 3: Retrieval-augmented generation (RAG) for evidence synthesis. For each sub-question $q_{i}$, we employ a Retrieval-Augmented Generation (RAG) framework to gather and synthesize relevant evidence: a Section retrieval: For each paper $P$, we retrieve pertinent text segments $S$ (approximately 400 tokens) via semantic search. b Parent section extraction: Using LangChain’s Parent Document Retriever, we extract the parent sections $S_{p}$ (approximately 4000 tokens) corresponding to each $S$. Documents are pre-chunked hierarchically, ensuring that each $S$ is mapped to its contextually relevant $S_{p}$. c Answer generation: With the context provided by $S_{p}$, the LLM generates an answer $A_{i}$ for each sub-question $q_{i}$. These individual answers are then aggregated into a unified, structured response $A_{Q}$ addressing the original question $Q$.
Step 4: Automated rebuttal generation. The comprehensive answer $A_{Q}$ is used to generate an automated rebuttal $R_{b}$ for the original review $R$. This rebuttal is designed to provide evidence-based clarification or counterarguments to the claims in $R$.
Step 5: Factual correctness evaluation. An evaluation agent then assesses the factual correctness of $R$ by comparing it against the generated rebuttal $R_{b}$. The review is deemed: a $\textbf{Valid}\ (\mathcal{V}=\text{True})$ if $R_{b}$ substantiates the claims made in $R$. b $\textbf{Invalid}\ (\mathcal{V}=\text{False})$ if $R_{b}$ reveals factual discrepancies or unsupported claims in $R$.

### v3

@@ Abstract
The escalating volume of academic research, coupled with a shortage of qualified reviewers, necessitates innovative approaches to peer review. In this work, we propose: ❶ ReviewEval, a comprehensive evaluation framework for AI-generated reviews that measures alignment with human assessments, verifies factual accuracy, assesses analytical depth, identifies degree of constructiveness and adherence to reviewer guidelines; and ❷ ReviewAgent, an LLM-based review generation agent featuring a novel alignment mechanism to tailor feedback to target conferences and journals, along with a self-refinement loop that iteratively optimizes its intermediate outputs and an external improvement loop using ReviewEval to improve upon the final reviews. ReviewAgent improves actionable insights by 6.78% and 47.62% over existing AI baselines and expert reviews respectively. Further, it boosts analytical depth by 3.97% and 12.73%, enhances adherence to guidelines by 10.11% and 47.26% respectively. This paper establishes essential metrics for AI-based peer review and substantially enhances the reliability and impact of AI-generated reviews in academic research.
@@
Madhav Krishan Garg1, Tejash Prasad1, Tanmay Singhal1, Chhavi Kirtani1, Murari Mandal2, Dhruv Kumar1, 3, 1IIIT Delhi, 2KIIT Bhubaneswar, 3BITS Pilani, Correspondence: dhruv.kumar@pilani.bits-pilani.ac.in
@@ 1 Introduction
The rapid growth of academic research, coupled with a shortage of qualified reviewers, has created an urgent need for scalable and high-quality peer review processes Petrescu and Krishen (2022); Schulz et al. (2022); Checco et al. (2021). This need has led to a growing interest in leveraging large language models (LLMs) to automate and enhance various aspects of the peer review process Robertson (2023); Liu and Shah (2023).
Although LLMs have shown remarkable potential in automating various natural language processing tasks, their effectiveness in serving as reliable and consistent paper reviewers remains a significant challenge. The academic community is already experimenting with AI-assisted reviews; for instance, 15.8% of ICLR 2024 reviews involved AI assistance Latona et al. (2024). Despite increasing LLM adoption, concerns about reliability and fairness remain.
Figure 1: ReviewEval and ReviewAgent: Given a paper and the associated conference or journal guidelines, ReviewAgent generates AI-based reviews and evaluates them along multiple dimensions using ReviewEval.
Specifically, papers reviewed by AI have been perceived to gain an unfair advantage, leading to questions about the integrity of such evaluations. Consequently, research into robust automated review generation systems is crucial, necessitating rigorous evaluation of AI generated reviews to address key challenges. Zhou et al. (2024) analyze commercial models like GPT-3.5 and GPT-4 Achiam et al. (2023) as paper reviewers, identifying limitations such as hallucinations, incomplete understanding, and insufficient critical feedback compared to humans.
@@
Existing research on evaluation metrics for AI-generated research paper reviews remains limited. For instance, D’Arcy et al. (2024) proposed a GPT-4-based automated metric for evaluating approximate matches between AI-generated and human-written reviews. However, their method’s iterative reliance on GPT-4 creates a black-box evaluation, limiting transparency and raising reliability concerns. Similarly, Zhou et al. (2024) investigated the aspect coverage and similarity between AI and human reviews through a blend of automatic metrics and manual analysis. But their approach overlooks other critical dimensions where AI reviews may underperform, as highlighted in Figure 2.
@@
Addressing the above mentioned limitations and gaps, we propose ReviewEval, a comprehensive evaluation framework for assessing the quality of AI-generated research paper reviews. ReviewEval targets five key dimensions (see Figure 2): ❶ Comparison with Human Reviews: Evaluates topic coverage and semantic similarity to measure the alignment between AI-generated and human-written feedback. ❷ Factual Accuracy: Detects factual errors, including misinterpretations, incorrect claims, and hallucinated information. ❸ Analytical Depth: Assesses whether the AI’s critique transcends generic commentary to offer in-depth, meaningful feedback. ❹ Actionable Insights: Measures the ability of the AI to provide constructive suggestions for improving the paper. ❺ Adherence to Reviewer guidelines: Quantifies the degree to which a review conforms to the evaluation criteria outlined by the target conference.
@@
Recent studies Bauchner and Rivara (2024); Biswas (2024) underscore the growing importance of aligning reviews with conference-specific evaluation criteria, especially as many venues now require adherence to detailed reporting guidelines. To address this, we introduce ReviewAgent, an AI reviewer with three key features: ❶ Conference-specific review alignment, dynamically adapting reviews to each venue’s unique evaluation criteria; ❷ Self-refinement loop, iteratively optimizing prompts and intermediate outputs for deeper analytical feedback; ❸ External improvement loop, systematically enhancing review quality using evaluation metrics provided by ReviewEval.
Using ReviewEval, we extensively evaluate ReviewAgent against two baseline models, AI-Scientist  Lu et al. (2024) and MARG  D’Arcy et al. (2024), benchmarking all systems against expert-written reviews treated as the gold standard. The experiments are performed on a balanced and topically diverse dataset of 120 papers sampled from NeurIPS, ICLR, and UAI, including domains such as computer science, biology, social science, finance, physics, and engineering.
Our results show that ReviewAgent improves actionable insights by 6.78% and 47.62% over existing AI baselines and expert reviews respectively. Further, it boosts analytical depth by 3.97% and 12.73%, enhances adherence to guidelines by 10.11% and 47.26% respectively.
In summary, the paper makes the following research contributions: (1) ReviewEval: a multi-dimensional framework for evaluating research paper reviews (§3). (2) ReviewAgent: an adaptive AI reviewer which utilizes ReviewEval along with novel conference alignment and iterative self-improvement (§4). (3) Comprehensive evaluation of ReviewAgent against state-of-the-art baselines using ReviewEval (§5).
@@ 2 Related Work
AI-based scientific discovery. Early attempts to automate research include AutoML that optimize hyperparameters and architectures Hutter et al. (2019); He et al. (2021) and AI-driven discovery in materials science and synthetic biology Merchant et al. (2023); Hayes et al. (2024). However, these methods remain largely dependent on human-defined search spaces and predefined evaluation metrics, limiting their potential for open-ended discovery. Recent works Lu et al. (2024) aim to automate the entire research cycle, encompassing ideation, experimentation, manuscript generation, and peer review, thus pushing the boundaries of AI-driven scientific discovery.
Figure 2: Illustrative examples highlighting key challenges in AI-generated research paper reviews that motivate our proposed evaluation metrics: Column 1 shows semantic and topical divergences from human reviews, supporting the Human Comparison metric; Column 2 presents factual inaccuracies or hallucinations, motivating the Factual Accuracy metric; Column 3 illustrates limited analytical reasoning, justifying the Analytical Depth metric; Column 4 reveals a lack of specific, actionable suggestions, underscoring the need for the Actionable Insights metric.
@@
Lu et al. (2024) employ LLMs to autonomously conduct the research pipeline, including peer review. It follows a structured three-stage review process: paper understanding, criterion-based evaluation (aligned with NeurIPS and ICLR guidelines), and final synthesis: assigning scores to key aspects like novelty, clarity, and significance. MARG D’Arcy et al. (2024) introduces a multi-agent framework where worker agents review sections, expert agents assess specific aspects, and a leader agent synthesizes feedback. Using BERTScore Zhang* et al. (2020) and GPT-4-based evaluation, MARG-S improves feedback quality, reducing generic comments and increasing helpful feedback per paper. These studies highlight the AI’s potential to enhance peer review through structured automation and multi-agent collaboration.
Evaluation framework for AI-based peer-review. There has been limited research on developing evaluation frameworks for evaluating the quality of LLM generated paper reviews. Zhou et al. (2024) evaluated GPT models for research paper reviewing across 3 tasks: aspect score prediction, review generation, and review-revision MCQ answering. Their evaluation framework comprised aspect coverage (originality, soundness, substance, replicability, etc.), ROUGE (lexical overlap), BERTScore (semantic similarity), and BLANC (informativeness), alongside manual analysis. Results showed LLMs overemphasized positive feedback, lacked critical depth, and neglected substance and clarity, despite high lexical similarity to human reviews. D’Arcy et al. (2024) introduced an automated evaluation framework for AI-generated reviews, quantifying similarity to human reviews via recall, precision, and Jaccard index. Recall measures the fraction of real-reviewer comments with at least one AI match, precision quantifies AI comments aligned with human reviews, and Jaccard index evaluates the intersection-over-union of aligned comments.
@@
Iterative Refinement of Large Language Models. Recent studies highlight the benefits of iterative feedback-driven refinement in improving LLM outputs: Self-Refine Madaan et al. (2023) uses self-generated feedback without additional training, LLMRefine Xu et al. (2024) employs a learned feedback model with simulated annealing, and ProMiSe Hu et al. (2024) leverages external proxy metrics, collectively enhancing factuality, coherence, and performance in tasks like document-grounded QA and dialog generation.
@@
Table 1: Existing metrics focus on AI-human review similarity but overlook other key aspects; our framework fills these gaps with more interpretable metrics.
@@ 3 ReviewEval
We introduce ReviewEval, an evaluation framework grounded in the LLM-as-a-Judge paradigm Gu et al. (2025). In ReviewEval, each review is evaluated on several key parameters to assess the overall quality of the generated feedback. To ensure consistency and reliability, all evaluations for a given metric were performed by LLMs of the same specification and version, minimizing variability from model differences and enabling robust, fair, and unbiased comparisons. The overview of the proposed framework (ReviewEval + ReviewAgent) is presented in Figure 1.
@@ 3.1 Comparison with Expert Reviews
We compare the reviews generated by the LLM based reviewer with expert human reviews. Our primary goal is to gauge how well the AI system replicate expert-level critique. The evaluation is conducted along the following dimensions:
Semantic similarity. To assess the alignment between AI-generated and expert reviews, we embed each review $R$ into a vector representations using a text embedding model Mikolov et al. (2013). The semantic similarity between an AI-generated review $R_{\text{AI}}$ and an expert review $R_{\text{Expt}}$ is measured using cosine similarity:
@@
Topic Coverage. We evaluate topic coverage to determine how comprehensively AI-generated reviews address the breadth of topics present in expert reviews. Our approach comprises three steps: ❶ Topic extraction: Each review $R$ (either AI-generated or expert) is decomposed into a set of topics by an LLM: $T_{R}=\{t_{1},t_{2},\dots,t_{n}\},$ where each topic $t_{i}$ is represented by a sentence that captures its core content and context. ❷ Topic similarity: Let $T_{\text{AI}}=\{t_{1},t_{2},\dots,t_{m}\}$ and $T_{\text{Expt}}=\{t^{\prime}_{1},t^{\prime}_{2},\dots,t^{\prime}_{n}\}$ denote the topics extracted from the AI and expert reviews, respectively. We define a topic similarity function $\text{TS}(t_{i},t^{\prime}_{j})$ that an LLM assigns on a discrete scale:
$\text{TS}(t_{i},t^{\prime}_{j})=3\cdot\mathbb{I}\{t_{i}\sim_{\text{strong}}t^{\prime}_{j}\}+2\cdot\mathbb{I}\{t_{i}\sim_{\text{moderate}}t^{\prime}_{j}\}+1\cdot\mathbb{I}\{t_{i}\sim_{\text{weak}}t^{\prime}_{j}\}$, where $\mathbb{I}$ is the indicator function, $t_{i}\sim_{\text{strong}}t^{\prime}_{j}$, $t_{i}\sim_{\text{moderate}}t^{\prime}_{j}$ , $t_{i}\sim_{\text{weak}}t^{\prime}_{j}$ denote substantial, moderate, and minimal overlap in concepts, respectively. All the conditions are mutually exclusive. We set a similarity threshold $\tau=2$ so that a topic in AI-generated review having atleast moderate similarity to a topic in expert review is considered aligned. ❸ Coverage ratio: For each AI-generated review, we construct a topic similarity matrix $S$ where each element $S[i,j]=\text{TS}(t_{i},t^{\prime}_{j})$ represents the similarity between topic $t_{i}$ from $T_{\text{AI}}$ and topic $t^{\prime}_{j}$ from $T_{\text{Expt}}$. The topic coverage ratio is defined as:
@@
3.2 Factual Correctness
To address hallucinations and factual inaccuracies in LLM-generated reviews, we introduce an automated pipeline that simulates the conference rebuttal process, allowing evidence-based validation of reviewer claims. By automating both the question-generation and rebuttal phases, our system produces a robust factual correctness evaluation. The pipeline consists of the following steps:
Step 1: Question generation. Each review $R$ is converted by an LLM into a structured verification question $Q$ that captures its central critique (Example in Appendix A.2).
Step 2: Query decomposition. A dedicated LLM-based decomposition component splits $Q$ into a set of sub-questions $\{q_{1},\ldots,q_{n}\}$, isolating distinct aspects of the claim for fine-grained analysis. (Implementation and Example in Appendix A.2.)
Step 3: Retrieval-augmented synthesis. For each sub-question $q_{i}$, we: a retrieve relevant text segments ($\approx 1000$ tokens) via semantic similarity search over contextual embeddings; b extract the corresponding parent sections ($\approx 4000$ tokens) to provide broader context. Both parent and child document splitters use a 10% overlap to enhance context retention across chunk boundaries; c generate a focused answer $A_{i}$ to $q_{i}$ using the retrieved context. The individual answers are then aggregated into a unified response $A_{Q}$ using an LLM addressing the original question $Q$.
Step 4: Automated rebuttal generation. The aggregated evidence $A_{Q}$ serves as the basis for an evidence-based rebuttal $R_{b}$, which systematically supports or counters each claim in $R$ by citing specific sections of the paper.
Step 5: Factual correctness evaluation. We compare $R$ against $R_{b}$ to determine whether each claim is substantiated:
$\mathcal{V}=\left\{\begin{array}[]{ll}1\!:&\text{all claims in }R\text{ are supported by }R_{b};\\ 0\!:&\text{otherwise}\end{array}\right.$
We then compute an overall factual correctness score:
$S_{\text{factual}}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\mathbb{I}\bigl{(}\mathcal{V}_{i}=\text{1}\bigr{)},$ (3)